alternative are mutually and have infinite useful lives determine which alternative should be selected for the investment using İNTERNAL RATE OF RETURN (IRR METHOD) (Marr %10)

project a project b project c

ınıtıal ınvestment cost 10.000.000 8.500.000 7.600.000

annual revenues 9.500.000 900.000 800.000

salvage value 5.000.000 1.000.000 0
Answer #1

Answer:

Project A Project B Project C
Initial Investment P 10,000,000 8,500,000 7,600,000
Annual Revenue A 9,500,000 900,000 800,000
Salvage Value S 5,000,000 1,000,000 0

for IRR calculation

P=A/IRR

IRR=A/P

PV of Salvage value will be zero for infinite useful life.

For Project A

IRR=9,500,000/10,000,000=95%

If Annual return for A is 950,000

Then IRR=950000/10,000,000=9.5%

For Project B

IRR=900,000/8,500,000=10.58%

For Project C

IRR=800,000/7,600,000=10.53%

If Annual return for A=9,500,000

Since all the project has IRR greater than MARR (10%) so project with highest IRR will be considered which project isA.

If Annual return for A=950,000

Since only two the project (B & C) has IRR greater than MARR (10%) so project with highest IRR will be considered which is project B.